OpenStudy (anonymous):

Given that : a=b -cd, work out the value of a when b=-1,c=-2 and d=3

OpenStudy (anonymous):

a=b-cd a=-1+2(3) a=5

OpenStudy (anonymous):

thxxxx

OpenStudy (anonymous):

Anytime :)